Find the equation of the line tangent to the circle with the center at (-1,1) and point of tangency at (-1,3).

1
Expert's answer
2021-12-09T07:43:05-0500

The equation of the circle


"(x-(-1))^2+(y-1)^2=R^2"

The point "(-1,3)" lies on the circle


"(-1-(-1))^2+(3-1)^2=R^2"

"R^2=4"

"(x+1)^2+(y-1)^2=4"

Differentiate both sides with respect to "x"


"2(x+1)+2(y-1)y'=0"

Solve for "y'"


"y'=-\dfrac{x+1}{y-1}"

Point of tangency at "(-1,3)."


"y'=-\dfrac{-1+1}{3-1}=0"

The equation of the tangent in point-slope form


"y-3=0(x-(-1))"

The equation of the tangent in slope-intercept form


"y=3"